Mathematics 242 – Principles of Analysis Definitions Quiz 3 – February 15, 2013

Your Name:

A) (5) Define: The sequence {x_{n}} converges to a ∈ R.

B) (5) Let S = {x^{2} − 4x | x ∈ [0, 4]}. What is lub(S)?